OpenStudy (mathmath333):

A can hit the target 3 times in 6 shots, B 2 times in 6 shots and C 4 times in 6 shots. They fire a volley. What is the probablity that at least 2 shots hit ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ A can hit the target 3 times in 6 shots, B 2 times in 6 shots }\hspace{.33em}\\~\\ & \normalsize \text{ and C 4 times in 6 shots. They fire a volley. What is the }\hspace{.33em}\\~\\ & \normalsize \text{ probablity that at least 2 shots hit ?}\hspace{.33em}\\~\\ & a.)\ \dfrac{1}{2} \hspace{.33em}\\~\\ & b.)\ \dfrac{1}{3} \hspace{.33em}\\~\\ & c.)\ \dfrac{2}{3} \hspace{.33em}\\~\\ & d.)\ \dfrac{3}{4} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

A can hit the target 3 times in 6 shots, B 2 times in 6 shots and C 4 times in 6 shots. They fire a volley. What is the probablity that at least 2 shots hit ? a.)1/2 b.)1/3 c.)2/3 d.)3/4

OpenStudy (badhi):

\[p = P(A \cap B) +P(B \cap C) +P(A \cap C) +P(A \cap B \cap C) \]

OpenStudy (mathmath333):

i calculated 5/6 by ur formula

OpenStudy (mathmath333):

is it correct

OpenStudy (jadedry):

So A, B and C will be fired all at once? Okay. prob A can hit the target: 1/2 prob B:1/3 prob C: 2/3 If at least 2 targets have to be hit, we should calculate the combined probability of the following possibilities: No targets are going to be hit. \[\frac{ 1 }{ 2 } *\frac{ 2 }{ 3 } * \frac{ 1 }{ 3 } = \frac{ 1 }{ 9 }\] 1 target will be hit. For this scenario, there are multiple permutations. A hits, but not B/C - prob = 1/9 B hits, but not A/C - prob = 1/18 C hits, but not A/B - prob = 2/9 added together, you have a probability of: 7/18 Add to the probability that none of them hit. 7/18 + 1/9 = 9/18 The answer should be 1/2

OpenStudy (jadedry):

I should add that you must subtract the probability "none of them hit" from 1 to get the right value, but since the value (according to my calculations) was 1/2, it would have been the same either way.

OpenStudy (anonymous):

I think its 1/3 because 2/6 = 1/3

OpenStudy (phi):

Jade's answer looks good to review, using \(\bar{A}\) to mean A missed Chance of hit and miss are as follows A 1/2 \(\bar{A}\) 1/2 B 1/3 \(\bar{B}\) 2/3 C 2/3 \(\bar{C}\) 1/3 one way to do it is Pr(2 hits or more) = 1 - Pr(0 hits) - Pr(1 hit) To find Pr(0 hits) we multiply the chance each misses, namely \(\bar{A}\bar{B}\bar{C}\) =1/2 * 2/3 * 1/3 = 2/18 To find the Pr(1 hit) we add up the prob of \(AB\bar{C} + A\bar{B}C + \bar{A}BC= 1/2*1/3*1/3 + 1/2*2/3*2/3+1/2*1/3*2/3\) which is 1/18+4/18+2/18= 7/18 so we have Pr(2 hits or more) = 1 - Pr(0 hits) - Pr(1 hit) Pr(2 hits or more) = 1 -2/18 - 7/18 = 1 - 1/2 = 1/2

OpenStudy (badhi):

yeah I ve made a stupid mistake, my equation has to be \[\begin{align} p &= P(A\cap B \cap \bar C) + P(A\cap C \cap \bar B) + P(B\cap C \cap \bar A) + P(A\cap B \cap C)\\ &=P(A)P(B)P(\bar C)+P(A)P(\bar B)P( C)+P(\bar A)P(B)P(C)+P(A)P(B)P( C) \\ &= \left(\frac 1 2 \times\frac 1 3\times \frac 1 3 \right)+\left(\frac 1 2 \times\frac 2 3\times \frac 2 3 \right)+\left(\frac 1 2 \times\frac 1 3\times \frac 2 3 \right)+\left(\frac 1 2 \times\frac 1 3\times \frac 2 3 \right) \\&=\frac 1 2 \end{align}\] sorry guys